x-y=10 and 2x+3y=0 solve for x and y

Answer by addingup(3677)

x-y = 10 multiply both sides times 3:
3x-3y = 30 Now add this equation to the other equation:
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2x+3y = 0
+
3x-3y = 30
-----------
5x = 30 Divide both sides by 5:
x = 6
Now that we know x, let's find y:
x-y = 10
6-y = 10 multiply both sides times -1
-6+y = -10 Add 6 on both sides:
y = -4
And now that we have both, x and y, let's check and see if we have the right answer:
6--4= 6+4 = 10 Correct
2(6)+3(-4)= 0; 12+-12=0; 12-12= 0 Again, we have the right answer.
J